Two lines L1: 2y-3x-6=0 ad l2: 3y+x-20=0 intersect at a point A. Find the coordinates of A
Answer by Alan3354(69443)

Two lines L1: 2y-3x-6=0 ad l2: 3y+x-20=0 intersect at a point A. Find the coordinates of A
---------------------
2y-3x-6=0
3y+x-20=0
---
Multiply the 2nd equation by 3
9y + 3x - 60 = 0
2y - 3x - 6 = 0
----------------------------------- Add
11y - 66 = 0
11y = 66
y = 6
=================
2y - 3x - 6 = 0
12 - 3x - 6 = 0
-3x + 6 = 0
x = 2
----> A(2,6)